# symbol/.style= draw=none, every to/.append style= edge node=node [sloped, allow upside down, auto=false]#$1$

### v1

@@ 8.1. Semiabelian varieties
The following theorem is one of the statements that are usually called the Mordell-Lang conjecture for semiabelian varieties.
@@
8.2. Connected Shimura varieties
In this subsection, we work with the distinguished category of connected pure Shimura varieties. Let us start by considering the case of products of modular curves and see how to recover Theorem 1.8 of [2] for $Y(1)^{n}_{\mathbb{C}}$.
Let $Z=Y(1)^{n}_{\mathbb{C}}$ for some positive integer $n$ and let $\Sigma$ be a $\Sigma$-orbit. Recall that, in this setting, $\Sigma$-special subvarieties of $Z$ are $\Sigma_{Z}$-point special by Proposition 6.5. Then, Theorem 6.2 of [23] implies $\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ for all $m\geq 0$. Therefore, also $\operatorname{\mathrm{WML}}(Z,\Sigma,m)$ holds for all $m\geq 0$.
@@ 8.2. Connected Shimura varieties
Now, Theorem 7.8 would imply Theorem 1.8 of [2] for $Y(1)_{\mathbb{C}}^{n}$ if we could find a class $\mathcal{S}$ of distinguished varieties as described in the hypotheses of Theorem 7.8 and such that $\operatorname{\mathrm{WML}}(X,\Sigma,m)$ holds for all $X$ in $\mathcal{S}$ and for all $m\geq 0$. Proposition 6.6 of [15] (WCA is proved in [25]) gives us such a class $\mathcal{S}$ and it follows from the above that $\operatorname{\mathrm{WML}}(X,\Sigma,m)$ holds for all $X$ in $\mathcal{S}$. So we recover Theorem 1.8 of [2].
For a general connected pure Shimura variety $Z$, it follows from the considerations of Subsection 4.2 that, for every structure of finite rank $\widetilde{\Sigma}\subset Z(\mathbb{C})$ as defined in [1], there exists a $\Sigma$-orbit $\Sigma$ such that $\operatorname{\mathrm{WML}}(Z_{\mathbb{C}},\Sigma,m)$ for all $m\in\mathbb{N}$ implies the Mordell-Lang conjecture for $Z_{\mathbb{C}}$ and $\widetilde{\Sigma}$ as formulated as Conjecture 3.5 in [1]. However, while the Mordell-Lang conjecture for $Z_{\mathbb{C}}$ and a structure of finite rank as formulated in [1] is known if $Z$ is of abelian type (see Corollary 3.10 in [1]), we do not know if the same is true for our version of the conjecture.
@@ 8.3. Powers of elliptic schemes
Let $\pi:\mathcal{E}\to Y(2)=\mathbb{A}^{1}_{\bar{\mathbb{Q}}}\backslash\{0,1\}$ denote the Legendre family of elliptic curves over $\bar{\mathbb{Q}}$, defined in ¯$\mathbb{P}^{2}_{\bar{\mathbb{Q}}}\times Y(2)$ by the equation $Y^{2}Z=X(X-Z)(X-\lambda Z)$, where $[X:Y:Z]$ are the projective coordinates on ¯$\mathbb{P}^{2}_{\bar{\mathbb{Q}}}$ and $\lambda$ is the affine coordinate on $Y(2)$. For $n\in\mathbb{N}$, we denote by $\mathcal{E}^{(n)}$ the $n$-th fibered power of $\mathcal{E}$ over $Y(2)$ and by $\mathcal{E}^{n}$ the $n$-th direct power of $\mathcal{E}$ over $\bar{\mathbb{Q}}$. By abuse of notation, we also use $\pi$ to denote the projections $\mathcal{E}^{(n)}\to Y(2)$ and $\mathcal{E}^{n}\to Y(2)^{n}$. All the $\mathcal{E}^{(n)}$, $\mathcal{E}^{n}$, and $Y(2)^{n}$ are connected mixed Shimura varieties of Kuga type over $\bar{\mathbb{Q}}$.
@@
We claim that, for $S$ equal to $\mathcal{E}^{(n)}_{K}$, $\mathcal{E}^{n}_{K}$, $Y(2)^{n}_{K}$, and $E_{K}^{n}$ respectively, the $\Sigma$-special subvarieties of $S$ can be explicitly described like this: first, a subvariety $Y$ of $Y(2)_{K}^{n}$ is $\Sigma$-special if and only if there exist $s,t\in\mathbb{Z}_{\geq 0}$ such that, after a permutation of the coordinates, $Y$ becomes equal to $Y_{1}\times\cdots\times Y_{s}\times\{(y_{1},\ldots,y_{t})\}$, where for every $i=1,\ldots,s$, $Y_{i}$ is a $1$-dimensional irreducible component of the preimage in $Y(2)_{K}^{m_{i}}$ under the canonical Shimura morphism $Y(2)_{K}^{m_{i}}\to Y(1)_{K}^{m_{i}}$ of the common zero locus in $Y(1)_{K}^{m_{i}}$ of a finite set of modular polynomials evaluated at pairs of distinct coordinates, $\sum_{i=1}^{s}{m_{i}}+t=n$, and for every $i=1,\ldots,t$, either $y_{i}\in Y(2)(K)$ is special or $\mathcal{E}_{y_{i}}$ is isogenous to $\mathcal{E}_{\pi(x_{j})}$ for some $j\in\{1,\ldots,r\}$. Second, a subvariety $Y$ of $\mathcal{E}_{K}^{n}$ is $\Sigma$-special if and only if $\pi(Y)$ is a $\Sigma$-special subvariety of $Y(2)_{K}^{n}$ and, after a permutation of the coordinates, $Y$ is equal to $Y_{0}\times\{(y_{1},\ldots,y_{s})\}$ where $Y_{0}\subseteq\mathcal{E}_{K}^{n-s}$ ($0\leq s\leq n$) and the geometric generic fiber of $Y_{0}\to\pi(Y_{0})$ is a union of irreducible components of algebraic subgroups while for all $i=1,\ldots,s$, there exist $N_{i}\in\mathbb{N}$ and homomorphisms of algebraic groups $\phi_{i,j}:\mathcal{E}_{\pi(x_{j})}\to\mathcal{E}_{\pi(y_{i})}$ ($j=1,\ldots,r$) such that $N_{i}y_{i}=\sum_{j=1}^{r}{\phi_{i,j}(x_{j})}$ (note that $\phi_{i,j}=0$ is allowed). Finally, a subvariety of $\mathcal{E}_{K}^{(n)}$ or of $E_{K}^{n}$ is $\Sigma$-special if and only if it is $\Sigma$-special when considered as a subvariety of $\mathcal{E}_{K}^{n}$ (where we identify $E$ with some fiber of $\mathcal{E}\to Y(2)$ that is isomorphic to it – the choice of the fiber does not matter).
Indeed, this claim follows from Lemma 3.7 together with the characterization of special subvarieties of products of modular curves in Propositions 2.1 and 3.1 in [9] and the characterization of special subvarieties of connected mixed Shimura varieties of Kuga type in Proposition 5.3 in [14] together with Proposition 5.13 in [13] and Proposition 3.4 and its proof in [12].
@@
We apply our above characterization of the $\Sigma$-special subvarieties of $\mathcal{E}^{(n)}_{K}$. We also use again Proposition 5.3 in [14], which yields a characterization of the weakly special subvarieties of $\mathcal{E}^{(n)}_{K}$. Under condition (1), $\Sigma$ is just the set of special points of $\mathcal{E}^{(n)}_{K}$ and Mordell-Lang reduces to André-Oort. In this case the claim follows from Theorem 13.6 in [13]. Under condition (2), we deduce the theorem by combining Theorem 13.6 in [13] with Theorem 1.1 in [7]. Indeed, given a subvariety $V$ of $\mathcal{E}^{(n)}_{K}$, if it contains a dense set of special points, then it is special by André-Oort. Otherwise, in $V$ there must be a dense set of non-special points $y$ that are $\Sigma$-special. This means that, for any such $y=(y_{1},\dots,y_{n})$, $\mathcal{E}_{\pi(y)}$ is isogenous to one of the $\mathcal{E}_{\pi(x_{k})}$ without CM and relations of the form $N_{i}y_{i}=\sum_{j=1}^{r}{\phi_{i,j}(x_{j})}$ hold, where the $N_{i}$ are positive integers and $\phi_{i,j}:\mathcal{E}_{\pi(x_{j})}\to\mathcal{E}_{\pi(y_{i})}$ ($j=1,\ldots,r$) are homomorphisms of algebraic groups ($i=1,\ldots,n$). Our claim follows from Theorem 1.1 in [7]. Under condition (3), if $K=\bar{\mathbb{Q}}$, the theorem follows from Theorem 1.3 in [8]. For general $K$, a sketch of the proof can be found in Section 5.1 in [6]. ∎
@@
Let $n\in\mathbb{N}$ and let $m$ be an non-negative integer. Suppose either that $m\leq 1$ or that $K$ and $x$ satisfy condition (1) or (2) in Theorem 8.2. Then, if $V\subset\mathcal{E}^{(n)}_{K}$ is a subvariety of dimension at most $m$, there are at most finitely many $\Sigma$-optimal varieties for $V$ in $\mathcal{E}^{(n)}_{K}$.
@@
We want to combine Theorems 7.8 and 8.2. We have to verify that $\operatorname{\mathrm{WML}}(X^{\prime},\Sigma,m)$ holds for every distinguished variety $X^{\prime}$ in a class $\mathcal{S}$ that verifies the hypothesis in Theorem 7.8 for $Z=\mathcal{E}^{(n)}_{K}$. Arguing as in the proof of Theorem 13.1 in [3], we see that it suffices to check that $\operatorname{\mathrm{WML}}(\mathcal{E}^{(k)}_{K},\Sigma,m)$ and $\operatorname{\mathrm{WML}}(E^{k}_{K},\Sigma,m)$ hold for all $k\in\mathbb{N}$ and all elliptic curves with complex multiplication $E$ over $\bar{\mathbb{Q}}$ (regarded as connected mixed Shimura varieties of Kuga type over $\bar{\mathbb{Q}}$). But this follows from Theorem 8.2. ∎

### v2

@@ 7. Mordell-Lang
Under an additional hypothesis that is often satisfied in the applications, the two statements are equivalent.
Lemma 7.3.
Let $\Sigma$ be a $\Sigma$-orbit in a distinguished category that satisfies (A5) and let $Z$ be a distinguished variety. If every $\Sigma$-special subvariety of $Z$ contains a dense set of $\Sigma$-special points, then $\operatorname{\mathrm{WML}}(Z,\Sigma,m)\Leftrightarrow\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ for all non-negative integers $m$.
Proof.
One implication has already been discussed above. For the other implication, let $V$ be a subvariety of $Z$ of dimension $\leq m$ and let $S\subset V$ be a maximal $\Sigma$-special subvariety. The Zariski closure of all $\Sigma$-special points in $V$ contains $S$ and is a finite union of weakly special and hence $\Sigma_{Z}$-point-special subvarieties by $\operatorname{\mathrm{WML}}(Z,\Sigma,m)$. By Lemma 5.10, each component of this Zariski closure is also $\Sigma$-special. It follows from the maximality of $S$ that $S$ is equal to such a component and so there are at most finitely many possibilities for $S$. ∎
@@
The set $\mathcal{S}$ in Theorem 7.9 is a nuisance. We now prove two lemmas that allow us to get rid of it at the cost of assuming that $\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ holds for all $m\geq 0$.
Lemma 7.12.
In a distinguished category satisfying (A5), the distinguished morphisms $(\phi,\psi)$ in (A5) can be chosen such that each $\psi$ is surjective.
Proof.
Given $\phi:Y_{\phi}\to X$ and $\psi:Y_{\phi}\to Z_{\psi}$ such that $\phi$ has finite fibers, we apply (A4) to find distinguished morphisms $\phi_{1}:V_{\phi}\to Y_{\phi}$, $\psi_{1}:V_{\phi}\to U_{\psi}$, and $\psi_{2}:U_{\psi}\to Z_{\psi}$ such that $\psi\circ\phi_{1}=\psi_{2}\circ\psi_{1}$, $\phi_{1}$ is finite and surjective, $\psi_{1}$ is surjective, and $\psi_{2}$ has finite fibers.
Let $W$ be an irreducible component of $\phi(\psi^{-1}(z))$ for some $z\in Z_{\psi}(K)$. Note that
$\phi(\psi^{-1}(z))=(\phi\circ\phi_{1})((\psi\circ\phi_{1})^{-1}(z))=(\phi\circ\phi_{1})((\psi_{2}\circ\psi_{1})^{-1}(z))=\bigcup_{u\in\psi_{2}^{-1}(z)}{(\phi\circ\phi_{1})(\psi_{1}^{-1}(u))},$
where the last union of closed subsets is finite since $\psi_{2}$ has finite fibers. Hence, $W$ is also an irreducible component of $(\phi\circ\phi_{1})(\psi_{1}^{-1}(u))$ for some $u\in U_{\psi}(K)$. Therefore, we can replace $(\phi,\psi)$ by $(\phi\circ\phi_{1},\psi_{1})$, where $\phi\circ\phi_{1}$ has finite fibers and $\psi_{1}$ is surjective. ∎
Lemma 7.13.
Let $\Sigma$ be a $\Sigma$-orbit and let $X$ be a distinguished variety such that, for all integers $m\geq 0$, $\operatorname{\mathrm{SML}}(X,\Sigma,m)$ holds. Let $\phi:Y\to X$ be a distinguished morphism with finite fibers and let $\psi:Y\to Z$ be a surjective distinguished morphism. Then $\operatorname{\mathrm{SML}}(Y,\Sigma,m)$ and $\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ hold for all integers $m\geq 0$.
Proof.
We first prove that $\operatorname{\mathrm{SML}}(Y,\Sigma,m)$ holds for all integers $m\geq 0$. We proceed by induction on $m$, the case $m=0$ being trivial. So let $V\subset Y$ be a subvariety of dimension $m$ and let $S\subset V$ be a maximal $\Sigma$-special subvariety. Let $W$ denote the closure of $\phi(V)$. Then, by Proposition 5.11, $\phi(S)$ is a $\Sigma$-special subvariety of $W$ and therefore contained in a maximal $\Sigma$-special subvariety $S^{\prime}$ of $W$, which must come from a finite set since $\operatorname{\mathrm{SML}}(X,\Sigma,m)$ holds for all $m$.
We have $S\subset\phi^{-1}(S^{\prime})$, so $S$ is contained in some component $S^{\prime\prime}$ of $\phi^{-1}(S^{\prime})$. We first assume that $S^{\prime\prime}\not\subset V$. Since $\phi$ has finite fibers, we have $\dim S^{\prime\prime}\leq\dim S^{\prime}\leq\dim W=\dim V$. Hence, we cannot have $V\subset S^{\prime\prime}$ because otherwise $S^{\prime\prime}=V$, a contradiction. So $S^{\prime\prime}\cap V$ is a proper closed subset of $V$ that contains $S$ and we are again done by induction (note that $S^{\prime}$ and therefore also $S^{\prime\prime}$ vary in finite sets respectively).
We can therefore assume that $S^{\prime\prime}\subset V$. It follows from the maximality of $S$ that $S=S^{\prime\prime}$ and so $S$ lies in a finite set and we are done.
We now prove that $\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ holds for all integers $m\geq 0$. So let $V\subset Z$ be a subvariety. If $S\subset V$ is a maximal $\Sigma$-special subvariety, then every component of $\psi^{-1}(S)$ is a maximal $\Sigma$-special subvariety of some component of $\psi^{-1}(V)$, which is enough to conclude. ∎
@@ 8.1. Semiabelian varieties
The following theorem is one formulation of the Mordell-Lang conjecture for semiabelian varieties.
@@
8.2. Connected pure Shimura varieties
In this subsection, we work with the distinguished category of connected pure Shimura varieties. Let us start by considering the case of products of modular curves and see how to recover Theorem 1.8 of [2] for $Y(1)^{n}_{\mathbb{C}}$ from our Theorem 7.9.
Let $Z=Y(1)^{n}_{\mathbb{C}}$ for some positive integer $n$ and let $\Sigma=\Sigma(x)$ be a $\Sigma$-orbit for $x\in Y(1)^{n_{0}}(\mathbb{C})$ for some positive integer $n_{0}$. Recall that, in this setting, $\Sigma$-special subvarieties of $Z$ are $\Sigma_{Z}$-point-special by Proposition 6.5. Recall furthermore that we determined $\Sigma_{Z}$ in Subsection 4.2. Then, Theorem 6.2 of [26] (see also Theorem 3 in [18] for an earlier result) implies $\operatorname{\mathrm{SML}}(Z,\Sigma,m)$ for all $m\geq 0$. Therefore, also $\operatorname{\mathrm{WML}}(Z,\Sigma,m)$ holds for all $m\geq 0$.
@@ 8.2. Connected pure Shimura varieties
Now, Theorem 7.9 would imply Theorem 1.8 of [2] for $Y(1)_{\mathbb{C}}^{n}$ if we could find a class $\mathcal{S}$ of distinguished varieties as described in the hypotheses of Theorem 7.9 and such that $\operatorname{\mathrm{WML}}(X,\Sigma,m)$ holds for all $X$ in $\mathcal{S}$ and for all $m\geq 0$.
We directly get such a class $\mathcal{S}$ from Lemmas 7.12 and 7.13 in the following way: we start with the set $\{Y(1)_{\mathbb{C}}^{n};n\in\mathbb{N}\}$ and iteratively add distinguished varieties that admit a distinguished morphism with finite fibers to a distuingished variety in our set or that admit a surjective distinguished morphism from a distinguished variety from our set.
We now consider a general connected pure Shimura variety $Z$. Suppose that $\Sigma=\Sigma(x)$ for a point $x\in S(\mathbb{C})$, where $S$ is a connected pure Shimura variety. In Theorem 1.4 of [33], Richard and Yafaev have proved that $\operatorname{\mathrm{WML}}(Z_{\mathbb{C}},\Sigma,m)$ holds for all connected pure Shimura varieties $Z$ of abelian type and all $m$ (recall that in Subsection 4.2, we showed the equivalence of the notions of hybrid orbit and $\Sigma$-orbit). This implies that also $\operatorname{\mathrm{SML}}(Z_{\mathbb{C}},\Sigma,m)$ holds for all connected pure Shimura varieties $Z$ of abelian type and all $m$ (see Lemma 3.4 in [1] and Lemma 7.3). We can again use Lemmas 7.12 and 7.13 to get a class $\mathcal{S}$ that contains all connected pure Shimura varieties of abelian type and satisfies the condition in Theorem 7.9. Therefore, we obtain the following corollary of Theorem 7.9 together with the work of Richard and Yafaev.
Corollary 8.2.
Suppose that $\Sigma=\Sigma(x)$ for a point $x\in S(\mathbb{C})$, where $S$ is a connected pure Shimura variety. Let $Z$ be a connected pure Shimura variety of abelian type and let $V\subset Z$ be a subvariety. Then there are at most finitely many $\Sigma$-optimal varieties for $V$ in $Z$.
Recall from Subsection 4.2 that for every structure of finite rank $\widetilde{\Sigma}\subset Z(\mathbb{C})$ as defined in [1], there is a $\Sigma$-orbit $\Sigma=\Sigma(x)$ such that $\widetilde{\Sigma}\subset\Sigma_{Z_{\mathbb{C}}}$ and actually one can take $x$ to lie in $Z^{n}(\mathbb{C})$ for some $n\in\mathbb{N}$. Because of Proposition 6.5 and Remark 7.8, we then recover Theorem 1.2 of [1].
@@ 8.3. Powers of elliptic schemes
Let $\pi:\mathcal{E}\to Y(2)=\mathbb{A}^{1}_{\bar{\mathbb{Q}}}\backslash\{0,1\}$ denote the Legendre family of elliptic curves over $\bar{\mathbb{Q}}$, defined in ¯$\mathbb{P}^{2}_{\bar{\mathbb{Q}}}\times Y(2)$ by the equation $Y^{2}Z=X(X-Z)(X-\lambda Z)$, where $[X:Y:Z]$ are the projective coordinates on ¯$\mathbb{P}^{2}_{\bar{\mathbb{Q}}}$ and $\lambda$ is the affine coordinate on $Y(2)$. For $n\in\mathbb{N}$, we denote by $\mathcal{E}^{(n)}$ the $n$-th fibered power of $\mathcal{E}$ over $Y(2)$ and by $\mathcal{E}^{n}$ the $n$-th direct power of $\mathcal{E}$ over $\bar{\mathbb{Q}}$. By abuse of notation, we also use $\pi$ to denote the projections $\mathcal{E}^{(n)}\to Y(2)$ and $\mathcal{E}^{n}\to Y(2)^{n}$. All the $\mathcal{E}^{(n)}$, $\mathcal{E}^{n}$, and $Y(2)^{n}$ are connected mixed Shimura varieties of Kuga type over $\bar{\mathbb{Q}}$ (see Section 13 of [4]).
@@
We claim that, for $S$ equal to $\mathcal{E}^{(n)}_{K}$, $\mathcal{E}^{n}_{K}$, $Y(2)^{n}_{K}$, and $E_{K}^{n}$ respectively, the $\Sigma$-special subvarieties of $S$ can be explicitly described like this: first, a subvariety $Y$ of $Y(2)_{K}^{n}$ is $\Sigma$-special if and only if there exist $s,t\in\mathbb{Z}_{\geq 0}$ such that, after a permutation of the coordinates, $Y$ becomes equal to $Y_{1}\times\cdots\times Y_{s}\times\{(y_{1},\ldots,y_{t})\}$, where for every $i=1,\ldots,s$, $Y_{i}$ is a $1$-dimensional irreducible component of the preimage in $Y(2)_{K}^{m_{i}}$ under the canonical Shimura morphism $Y(2)_{K}^{m_{i}}\to Y(1)_{K}^{m_{i}}$ of the common zero locus in $Y(1)_{K}^{m_{i}}$ of a finite set of modular polynomials evaluated at pairs of distinct coordinates, $\sum_{i=1}^{s}{m_{i}}+t=n$, and for every $i=1,\ldots,t$, either $y_{i}\in Y(2)(K)$ is special or $\mathcal{E}_{y_{i}}$ is isogenous to $\mathcal{E}_{\pi(x_{j})}$ for some $j\in\{1,\ldots,r\}$. Second, a subvariety $Z$ of $\mathcal{E}_{K}^{n}$ is $\Sigma$-special if and only if $\pi(Z)$ is a $\Sigma$-special subvariety of $Y(2)_{K}^{n}$ and, after a permutation of the coordinates, $Z$ is equal to $Z_{0}\times Z_{1}$ where $Z_{0}\subseteq\mathcal{E}_{K}^{n-s}$ and $Z_{1}\subseteq\mathcal{E}^{s}_{K}$ ($0\leq s\leq n$) and the geometric generic fiber of $Z_{0}\to\pi(Z_{0})$ is a union of irreducible components of algebraic subgroups while $\pi(Z_{1})=\{(y_{1},\ldots,y_{s})\}$ for some $y_{1},\ldots,y_{s}\in Y(2)(K)$ and $Z_{1}$ is a translate of an abelian subvariety of $\mathcal{E}_{y_{1}}\times\cdots\times\mathcal{E}_{y_{s}}$ by a point $(z_{1},\ldots,z_{s})$ such that for all $i=1,\ldots,s$, there exist $N_{i}\in\mathbb{N}$ and homomorphisms of algebraic groups $\phi_{i,j}:\mathcal{E}_{\pi(x_{j})}\to\mathcal{E}_{y_{i}}$ ($j=1,\ldots,r$) such that $N_{i}z_{i}=\sum_{j=1}^{r}{\phi_{i,j}(x_{j})}$ (note that $\phi_{i,j}=0$ is allowed). Finally, a subvariety of $\mathcal{E}_{K}^{(n)}$ or of $E_{K}^{n}$ is $\Sigma$-special if and only if it is $\Sigma$-special when considered as a subvariety of $\mathcal{E}_{K}^{n}$ (where we identify $E$ with some fiber of $\mathcal{E}\to Y(2)$ that is isomorphic to it – the choice of the fiber does not matter).
Indeed, this claim follows from Lemma 5.4 together with the characterization of special subvarieties of products of modular curves in Propositions 2.1 and 3.1 in [11] and the characterization of Shimura subdata of connected mixed Shimura data of Kuga type in Proposition 1.2.16 in [14].
@@
We apply our above characterization of the $\Sigma$-special subvarieties of $\mathcal{E}^{(n)}_{K}$. We also use again Proposition 5.3 in [17], which yields a characterization of the weakly special subvarieties of $\mathcal{E}^{(n)}_{K}$. Under condition (1), $\Sigma$ is just the set of special points of $\mathcal{E}^{(n)}_{K}$ and Mordell-Lang reduces to André-Oort. In this case the claim follows from Theorem 13.6 in [16]. Under condition (2), we deduce the theorem by combining Theorem 13.6 in [16] with Theorem 1.1 in [9]. Indeed, given a subvariety $V$ of $\mathcal{E}^{(n)}_{K}$, if it contains a dense set of special points, then it is special by André-Oort. Otherwise, in $V$ there must be a dense set of non-special points $y$ that are $\Sigma$-special. This means that, for any such $y=(y_{1},\dots,y_{n})$, $\mathcal{E}_{\pi(y)}$ is isogenous to one of the $\mathcal{E}_{\pi(x_{k})}$ without CM and relations of the form $N_{i}y_{i}=\sum_{j=1}^{r}{\phi_{i,j}(x_{j})}$ hold, where the $N_{i}$ are positive integers and $\phi_{i,j}:\mathcal{E}_{\pi(x_{j})}\to\mathcal{E}_{\pi(y_{i})}$ ($j=1,\ldots,r$) are homomorphisms of algebraic groups ($i=1,\ldots,n$). Our claim then follows from Theorem 1.1 in [9].
Under condition (3), if $K=\bar{\mathbb{Q}}$, the theorem follows from Theorem 13.6 in [16] and Theorem 1.3 in [10]. For general $K$, a sketch of the proof of an analogue of Theorem 1.3 in [10] can be found in Section 5.1 in [8]. ∎
Remark 8.4.
See also Theorem 1.6 in [15] for a related result and see the master’s thesis [5] of Néstor Blázquez for a detailed characterization of the weakly special subvarieties of $\mathcal{E}^{(n_{1})}_{\mathbb{C}}\times\mathcal{E}^{(n_{2})}_{\mathbb{C}}$ as well as a proof of Theorem 8.3 under condition (2) without the assumption that $K=\bar{\mathbb{Q}}$.
@@
Let $n\in\mathbb{N}$. Suppose that $K$ and $x$ satisfy condition (1) or (2) in Theorem 8.3. Then, for every subvariety $V\subset\mathcal{E}^{(n)}_{K}$, there are at most finitely many $\Sigma$-optimal varieties for $V$ in $\mathcal{E}^{(n)}_{K}$.
Corollary 8.5 also holds if $V$ is a curve so that condition (3) from Theorem 8.3 is satisfied, but in that case, the conclusion of Corollary 8.5 is equivalent to $\operatorname{\mathrm{WML}}(\mathcal{E}^{(n)}_{K},\Sigma,1)$.
@@
We want to combine Theorems 7.9 and 8.3. We have to verify that $\operatorname{\mathrm{WML}}(X^{\prime},\Sigma,m)$ holds for all $m\geq 0$ and every distinguished variety $X^{\prime}$ in a class $\mathcal{S}$ that verifies the hypothesis in Theorem 7.9 for $Z=\mathcal{E}^{(n)}_{K}$. By Lemmas 7.12 and 7.13, it suffices to verify that $\operatorname{\mathrm{SML}}(\mathcal{E}^{(n)}_{K},\Sigma,m)$ holds for all $m\geq 0$ and $n\in\mathbb{N}$. Since every $\Sigma$-special subvariety of $\mathcal{E}^{(n)}_{K}$ contains a Zariski dense set of $\Sigma$-special points by the classification above, this follows from Lemma 7.3 and Theorem 8.3. ∎